# MLaKE: Multilingual Knowledge Editing Benchmark for Large Language Models

### v2

@@ Abstract
The extensive utilization of large language models (LLMs) underscores the crucial necessity for precise and contemporary knowledge embedded within their intrinsic parameters. Existing research on knowledge editing primarily concentrates on monolingual scenarios, neglecting the complexities presented by multilingual contexts and multi-hop reasoning. To address these challenges, our study introduces MLaKE (Multilingual Language Knowledge Editing), a novel benchmark comprising 4072 multi-hop and 5360 single-hop questions designed to evaluate the adaptability of knowledge editing methods across five languages: English, Chinese, Japanese, French, and German. MLaKE aggregates fact chains from Wikipedia across languages and utilizes LLMs to generate questions and answer. We assessed the effectiveness of current multilingual knowledge editing methods using the MLaKE dataset. Our results show that due to considerable inconsistencies in both multilingual performance and encoding efficiency, these methods struggle to generalize effectively across languages. The accuracy of these methods when editing English is notably higher than for other languages. The experimental results further demonstrate that models encode knowledge and generation capabilities for different languages using distinct parameters, leading to poor cross-lingual transfer performance in current methods. Transfer performance is notably better within the same language family compared to across different families. These findings emphasize the urgent need to improve multilingual knowledge editing methods. [^1] Our benchmark and source code are available at https://anonymous.4open.science/r/MLaKE_an-4D96/.
@@ 1 Introduction
To address this challenge, our study introduces MLaKE (Multi-lingual Language Knowledge Editing), a novel benchmark for multilingual multi-hop knowledge editing. MLaKE comprises 5360 single-hop questions and 4072 multi-hop questions designed to test the adaptability of knowledge editing methods across various languages. To ensure the quality and currency of knowledge, we begin by collecting fact chains across languages from Wikipedia. Subsequently, we leverage powerful LLM (e.g., ChatGPT) to generate questions in both free-form and multiple-choice formats using fact chains as input. Consequently, The MLaKE dataset comprises single-hop and multi-hop questions in English, Chinese, Japanese, French, and German. This diverse dataset serves as a robust foundation for evaluating the effectiveness of knowledge editing techniques in diverse linguistic environments.
@@ 2 MLaKE: Multi-Lingual Knowledge Editing Benchmark
In our study, we construct the MLaKE (Multi-Lingual Knowledge Editing) benchmark, an comprehensive and challenging dataset that encompasses five languages (English, Chinese, Japanese, French, German) and intricate logical structures (single-hop and multi-hop). In this section, we first present the data construction process of MLaKE, followed by a detailed description of the dataset. Lastly, we elaborate on the evaluation settings and metrics utilized.
@@ 2.1.3 Construction of question and answer
We employ ChatGPT to generate coherent and fluent multilingual questions, along with corresponding answers, from the collected triplet data in various languages. Instead of using the commonly applied template-based approach from previous research to convert triplets into questions, we opt for a different method due to the complexity of word inflections in French and German sentence structures. These complexities make natural conversion difficult when relying solely on template filling. To minimize potential errors in this transformation process, five experts in the relevant languages were invited to review the model-generated texts, following the criteria detailed in Appendix C.1.
@@ 2.2 Description of the MLaKE
Table 5 summarizes the statistics of the MLaKE dataset. The MLaKE dataset consists of more than 13K samples. We align all single-hop problems and employ them as a means to investigate the generalizability of existing knowledge editing methods across different languages following editing in a single language. Multi-hop problems are not aligned across different languages, and we employ them to assess the generalizability and portability of existing editing methods.
@@ 2.3 Evaluation Metrics
Diverging from other benchmarks, MLaKE primarily focuses on assessing the generalizability and portability of knowledge editing methods across multilingual scenarios. For different models, we use corresponding question to guide them to generate answers. We evaluate the accuracy of the question-answering task by determining whether the model-generated sentences contained the correct answer or its aliases.
@@ 3.1 Experimental Setup
  • Multilingual generalization refers to the accuracy of knowledge editing across different languages, especially the editing accuracy in other languages after using a single language for editing.
@@ 3.2 Generalization of Multilingual Knowledge Editing
To evaluate the multilingual generalization performance of commonly used knowledge editing methods. We select four widely used knowledge editing methods and evaluated their robustness across five different languages using the Vicuna and Qwen models. Each method are both edited and tested within the same language context. Table 1 presents the experimental results, and the main conclusion are as follows.
@@ 3.3 Transferability of Cross-Language Knowledge Editing
Conclusion 3: All knowledge editing methods have limited cross-language transferability, and the accuracy of the model in answering questions drops significantly when using different languages for reasoning. As illustrated in Figures 3, the editing performance along the diagonal is notably superior, indicating that the cross-language generalization ability of MEMIT and StableKE is limited. Notably, when performing knowledge editing on English on the Vicuna model, the accuracy in other languages was significantly lower than in English. A similar pattern was observed in the Qwen model during both Chinese and English editing.
Conclusion 4: The greater the similarity between the editing language and the reasoning language, the stronger the cross-language generalization ability. According to linguistic classification, English, German, and French all belong to the Indo-European language family Joseph (2005). Although Chinese and Japanese do not belong to the same language family, Japanese has been significantly influenced by Chinese Handel (2008); Jin-lian (2004). The figure reflects a trend: in the top-left corner, the cross-linguistic generalization among the three Indo-European languages is notably stronger, whereas generalization from these languages to Chinese and Japanese is significantly weaker than within the Indo-European family. While there is some connection between Chinese and Japanese, it is relatively weak, resulting in stronger cross-linguistic generalization between these two languages compared to their generalization to other languages, yet weaker than the generalization observed within the Indo-European language family.
@@ 4 Causes of Multilingual Editing Challenges
We identify two causes contributing to the suboptimal performance of multilingual knowledge editing: significant disparities in the multilingual capabilities of models and the structural independence of multilingual representations. The inconsistency in multilingual performance refers to the model’s varying proficiency and encoding efficiency across different languages, which undermines the generalization of knowledge editing in multilingual contexts. Structural independence in multilingual models suggests that knowledge and abilities in different languages are encoded by distinct parameters. This separation limits the effectiveness of knowledge editing in achieving cross-linguistic transfer.
@@ 5 Conclusion
In this paper, we explore the effectiveness of current knowledge editing methods in multilingual settings. To this end, we create the MLaKE dataset by extracting knowledge tuples from Wikipedia and generating single-hop and multi-hop questions using ChatGPT. Leveraging MLaKE, we conducte experiments employing various methods and multilingual LLMs to investigate the transferability of knowledge editing from English to other languages and vice versa. Our research findings reveal that: (1) Current knowledge-editing methods have limited generalization performance in multilingual settings. Beyond the inconsistent model performance across different languages, experiments suggest that insufficient multilingual encoding may contribute to their weak generalization. (2) The cross-lingual transferability of knowledge-editing methods is similarly constrained. Even within the same language family, transferability remains limited and is especially weak across distinct language families. A key finding from the experiments is that LLMs encode knowledge in different languages using distinct parameter sets, which reduces cross-lingual transfer performance.

### v3

@@ Abstract
The extensive utilization of large language models (LLMs) underscores the crucial necessity for precise and contemporary knowledge embedded within their intrinsic parameters. Existing research on knowledge editing primarily concentrates on monolingual scenarios, neglecting the complexities presented by multilingual contexts and multi-hop reasoning. To address these challenges, our study introduces MLaKE (Multilingual Knowledge Editing), a novel benchmark comprising 4072 multi-hop and 5360 single-hop questions designed to evaluate the adaptability of knowledge editing methods across five languages: English, Chinese, Japanese, French, and German. MLaKE aggregates fact chains from Wikipedia across languages and utilizes LLMs to generate questions and answer. We assessed the effectiveness of current multilingual knowledge editing methods using the MLaKE dataset. Our results show that due to considerable inconsistencies in both multilingual performance and encoding efficiency, these methods struggle to generalize effectively across languages. The accuracy of these methods when editing English is notably higher than for other languages. The experimental results further demonstrate that models encode knowledge and generation capabilities for different languages using distinct parameters, leading to poor cross-lingual transfer performance in current methods. Transfer performance is notably better within the same language family compared to across different families. These findings emphasize the urgent need to improve multilingual knowledge editing methods. [^1] Our benchmark and source code are available at https://github.com/Hi-archers/MLaKE.
@@ 1 Introduction
To address this challenge, our study introduces MLaKE (Multilingual Language Knowledge Editing), a novel benchmark for multilingual multi-hop knowledge editing. MLaKE comprises 5360 single-hop questions and 4072 multi-hop questions designed to test the adaptability of knowledge editing methods across various languages. To ensure the quality and currency of knowledge, we begin by collecting fact chains across languages from Wikipedia. Subsequently, we leverage powerful LLM (e.g., ChatGPT) to generate questions in both free-form and multiple-choice formats using fact chains as input. Consequently, The MLaKE dataset comprises single-hop and multi-hop questions in English, Chinese, Japanese, French, and German. This diverse dataset serves as a robust foundation for evaluating the effectiveness of knowledge editing techniques in diverse linguistic environments.
@@ 2 MLaKE: Multi-Lingual Knowledge Editing Benchmark
In our study, we construct the MLaKE (MultiLingual Knowledge Editing) benchmark, an comprehensive and challenging dataset that encompasses five languages (English, Chinese, Japanese, French, German) and intricate logical structures (single-hop and multi-hop). In this section, we first present the data construction process of MLaKE, followed by a detailed description of the dataset. Lastly, we elaborate on the evaluation settings and metrics utilized.
@@ 2.1.3 Construction of question and answer
Considering the complexities of inflection in French and German sentence structures, we avoid the template-driven methods often used in prior studies to convert triples into questions. Instead, we utilize ChatGPT to generate fluent and coherent multilingual questions, along with their corresponding answers, based on the collected triplet data in several languages. To minimize potential errors in this transformation process, five experts in the relevant languages were invited to review the model-generated texts, following the criteria detailed in Appendix C.1.
@@ 2.2 Description of the MLaKE
Table 5 summarizes the statistics of the MLaKE dataset. The MLaKE dataset consists of more than 13K samples. We align all single-hop problems and employ them as a means to investigate the generalizability of existing knowledge editing methods across different languages following editing in a single language. Multi-hop problems are not aligned across languages, and we use them to assess the generalizability and transferability of existing editing methods.
@@ 2.3 Evaluation Metrics
Diverging from other benchmarks, MLaKE primarily focuses on assessing the generalizability and transferability of knowledge editing methods across multilingual scenarios. For different models, we use corresponding question to guide them to generate answers. We evaluate the accuracy of the question-answering task by determining whether the model-generated sentences contained the correct answer or its aliases.
@@ 3.1 Experimental Setup
  • Multilingual generalization refers to the accuracy of knowledge editing across different languages, especially when edited knowledge is tested using single language.
@@ 3.2 Generalization of Multilingual Knowledge Editing
To evaluate the multilingual generalization performance of commonly used knowledge editing methods, we select four widely used knowledge editing methods and evaluated their robustness across five different languages using the Vicuna and Qwen models. Each method are both edited and tested within the same language context. Table 1 presents the experimental results, and the main conclusion are as follows.
@@ 3.3 Transferability of Cross-Language Knowledge Editing
Conclusion 3: All knowledge editing methods have limited cross-language transferability, and the accuracy of the model in answering questions drops significantly when using different languages for reasoning. As illustrated in Figures 3, the editing performance along the diagonal is notably superior, indicating that the cross-language transferability ability of MEMIT and StableKE is limited. Notably, when performing knowledge editing in English on the Vicuna model, the accuracy in other languages was significantly lower than in English. A similar pattern was observed in the Qwen model during both Chinese and English editing.
Conclusion 4: The greater the similarity between the editing language and the reasoning language, the stronger the cross-language transferability. According to linguistic classification, English, German, and French all belong to the Indo-European language family Joseph (2005). Although Chinese and Japanese do not belong to the same language family, Japanese has been significantly influenced by Chinese Handel (2008); Jin-lian (2004). The figure illustrates a trend: in the top-left corner, cross-linguistic transfer among the three Indo-European languages is notably stronger, whereas the transfer from these languages to Chinese and Japanese is significantly weaker than within the Indo-European family. While there is some linguistic connection between Chinese and Japanese, it is weaker compared to the Indo-European languages, resulting in stronger transfer between Chinese and Japanese than their transfer to other languages, but still weaker than the transfer observed within the Indo-European language family.
@@ 4 Causes of Multilingual Editing Challenges
We identify two causes contributing to the suboptimal performance of multilingual knowledge editing: significant disparities in the multilingual capabilities of models and the structural independence of multilingual representations. The inconsistency in multilingual performance refers to the model’s varying capability and encoding efficiency across different languages, which undermines the generalization of knowledge editing in multilingual contexts. Structural independence in multilingual models suggests that knowledge and abilities in different languages are encoded by distinct parameters. This separation limits the effectiveness of knowledge editing in achieving cross-linguistic transfer.
@@ 5 Conclusion
In this paper, we explore the effectiveness of current knowledge editing methods in multilingual settings. To this end, we create the MLaKE dataset by extracting knowledge tuples from Wikipedia and generating single-hop and multi-hop questions using ChatGPT. Leveraging MLaKE, we conduct experiments employing various methods and multilingual LLMs to investigate the generalization and transferability of knowledge editing from English and other languages. Our research findings reveal that: (1) Current knowledge-editing methods have limited generalization performance in multilingual settings. Beyond the inconsistent model performance across different languages, experiments suggest that insufficient multilingual encoding may contribute to their weak generalization. (2) The cross-lingual transferability of knowledge-editing methods is similarly constrained. Even within the same language family, transferability remains limited and is especially weak across distinct language families. A key finding from the experiments is that LLMs encode knowledge in different languages using distinct parameter sets, which reduces cross-lingual transfer performance.
@@
Acknowledgements
This work was supported by the Strategic Priority Research Program of the CAS under Grants No.XDB0680302, the National Natural Science Foundation of China (NSFC) under Grants No. 62276248, and the Youth Innovation Promotion Association CAS under Grants No. 2023111.